# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

### v1

@@ 5 Related works
Previous works on AAVE studies in natural language processing mostly focus on non-reasoning-heavy tasks such as POS tagging (Jørgensen et al., 2015; 2016), language identification and dependency parsing (Blodgett et al., 2016), automatic captioning (Tatman, 2017), and general language generation (Deas et al., 2023). AAVE is also found to be more likely to trigger false positives in hate speech identifiers (Davidson et al., 2019; Sap et al., 2019) due to specific word choices (Harris et al., 2022), be considered negative by automatic sentiment classifier (Groenwold et al., 2020), and cause covert biases in essential areas of social justice (Hofmann et al., 2024). Relevant studies (Ziems et al., 2022; Gupta et al., 2024) also find that rule-based AAVE feature perturbations can downgrade language model performance in a wide range of tasks covered by GLUE (Wang, 2018).
More generally, dialects in world languages pose challenges to natural language processing systems. Ziems et al. (2023) find that auto-encoder models are brittle on rule-based English dialect feature perturbations. Fleisig et al. (2024) report that English dialect speakers perceive responses generated by chatbots to be more negative than Standardized English prompts. Faisal et al. (2024) find that world dialects cause problems in tasks including dependency parsing (Scherrer et al., 2019) and machine translation (Mirzakhalov, 2021) on mBERT and XLM-R (Conneau et al., 2020).
However, existing works fail to systematically cover reasoning tasks in dialects. There is no existing high-quality end-to-end human-annotated dataset on such a task. Moreover, studies on LLM task-specific capabilities tend to focus on traditional auto-encoder models such as BERT and RoBERTa instead of SotA auto-regressive LLMs. Our work fills the gaps in these areas.
Fairness and Robustness of Large Language Models. LLMs are widely testified to be both unfair and brittle. They introduce unfair performance (Huang et al., 2023; Dong et al., 2024) and cost (Petrov et al., 2024) to users across different languages, exacerbate social imbalance by marginalizing minority groups in various aspects including gender (Kotek et al., 2023; Fraser & Kiritchenko, 2024), race (Hofmann et al., 2024; Wang et al., 2024), and culture (Naous et al., 2023; Tao et al., 2024). Our work shows for the first time that LLMs also exhibit unfairness in reasoning tasks for speakers of a dialect.
Previous works report that LLMs are very brittle to slight variations of prompts by introducing typos or paraphrasing in Standardized English (Elazar et al., 2021; Liang et al., 2022; Raj et al., 2022; Zhu et al., 2023b; Lin et al., 2024). In this work, we consider a novel application of using human-written perturbations in AAVE by asking humans to rewrite instances to their dialect and evaluate LLM robustness towards these natural perturbations, which have proven to cause LLMs to be more brittle than synthetic typo-style (Section 4.1) or linguistic-rule-based (Appendix A.8) perturbations.
@@ 6 Conclusion
In this work, we present ReDial which has $1.2K+$ parallel Standardized English-AAVE prompts to evaluate LLMs’ dialect robustness and fairness in algorithm, logic, math, and comprehensive reasoning as four canonical reasoning tasks. With ReDial, we find that SotA LLMs show significant unfairness and brittleness to reasoning tasks expressed in AAVE. The data skewness of AAVE does not explain the whole picture as large-scale LLMs are more brittle to AAVE compared to Standardized English typos of even higher perplexities. Prompting LLMs to rephrase questions in Standardized English cannot fully bridge the gap but tends to introduce higher costs. These findings highlight the unfairness of LLMs to dialect users and also shed light on the brittleness of LLMs’ reasoning capabilities when it comes to minor variations of prompts without changing their semantics. We call for further studies to enhance LLMs’ fairness and robustness to dialects to provide equal service to users from all linguistic groups and demographics.
7 Ethics Statement
@@
For data collection, we deliberately do not set hard constraints for annotator identity and demographic verification, recognizing there are no definite boundaries to identify dialects and their speakers (King, 2020) (we also note that the boundary between dialects and languages is sometimes difficult to define). King (2020) further elaborate that the term “AAVE” itself is contested, with alternatives that could be used instead; in employing the term “AAVE”, we adhere to the widely used terminology in related works on dialects and NLP (Ziems et al., 2022; Gupta et al., 2024). We corroborate the data quality by asking self-identified dialect speakers to cross-validate each others’ answers.
@@
Last, there are limitations on how well standard benchmarks reflect use cases of practical usage for LLMs. For ReDial, we select the source datasets among those reported in highly impactful LLM technical reports such as GPT-4 (Achiam et al., 2023), LLaMA-3 (Dubey et al., 2024), and Phi-3 (Abdin et al., 2024). Their popularity makes it easy to integrate them with existing pipelines, and the presence of ground truth labels mitigates inherent biases of using LLMs as evaluators (Zheng et al., 2023; Chen et al., 2024; Shi et al., 2024). Although we try our best to simulate user queries (e.g., changing code completion to instruction following queries in HumanEval), we do note there can be a gap between tasks as in standard benchmarks and queries in real workflows.
@@ 8 Acknowledgement
9 Reproducibility Statement
Code can be found at https://github.com/fangru-lin/redial_dialect_robustness_fairness.git and the dataset will be released upon publication.
@@ A.1.1 Algorithm
from typing import List def has_close_elements(numbers: List[float], threshold: float) -> bool: """Checkifingivenlistofnumbers,areanytwonumbers ␣␣␣␣closertoeachotherthangiventhreshold. ␣␣␣␣$>$has_close_elements([1.0,2.0,3.0],0.5) ␣␣␣␣False ␣␣␣␣$>$has_close_elements([1.0,2.8,3.0,4.0,5.0,2.0],0.3) ␣␣␣␣True ␣␣␣␣""" InstructHumanEval Used in the Paper
@@
Write a python function to remove first and last occurrence of a given character from the string. Your code should pass these tests: assert remove_Occ(“hello”,“l”) == “heo” assert remove_Occ(“abcda”,“a”) == “bcd” assert remove_Occ(“PHP”,“P”) == “H”
@@ A.1.2 Logic
Consider the following premises: “People in this club who perform in school talent shows often attend and are very engaged with school events. People in this club either perform in school talent shows often or are inactive and disinterested community members. People in this club who chaperone high school dances are not students who attend the school. All people in this club who are inactive and disinterested members of their community chaperone high school dances. All young children and teenagers in this club who wish to further their academic careers and educational opportunities are students who attend the school. Bonnie is in this club and she either both attends and is very engaged with school events and is a student who attends the school or is not someone who both attends and is very engaged with school events and is not a student who attends the school.” Assuming no other commonsense or world knowledge, is the sentence “Bonnie performs in school talent shows often.” necessarily true, necessarily false, or neither? Answer either “necessarily true”, “necessarily false”, or “neither”.
@@ A.4 Data Quality Verification
You will be given two prompts, one in Standard English and one in African American English. Determine whether the African American English prompt is a valid paraphrase of the Standard English prompt. Ignore the semantic validaty of the Standard English prompt. Standard English: ”[SAE_PROMPT]” African American English: ”[AAVE_PROMPT]” Is the African American English prompt a valid paraphrase of the Standard English prompt?
@@ A.7 Full Results on ReDial
Model Setting Original AAVE
[…]
CoT 0.762 0.662(-)0.1
[…]
CoT 0.283 0.058(-)0.225
[…]
CoT 0.075 0.067(-)0.008
LLaMA-3.1-70B-Instruct Vanilla 0.392 0.113(-)0.279
CoT 0.579 0.500(-)0.079
LLaMA-3-70B-Instruct Vanilla 0.158 0.067(-)0.092
CoT 0.517 0.350(-)0.167
LLaMA-3-8B-Instruct Vanilla 0.025 0.067(+)0.042
CoT 0.029 0.025(-)0.004
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.100 0.075(-)0.025
CoT 0.133 0.071(-)0.062
Mistral-7B-Instruct-v0.3 Vanilla 0.096 0.075(-)0.021
CoT 0.083 0.083(-)-0.0
Phi-3-Medium-128K-Instruct Vanilla 0.050 0.037(-)0.013
CoT 0.067 0.029(-)0.037
Phi-3-Small-128K-Instruct Vanilla 0.058 0.062(+)0.004
CoT 0.096 0.079(-)0.017
Phi-3-Mini-128K-Instruct Vanilla 0.021 0.042(+)0.021
CoT 0.017 0.021(+)0.004
Table 6: All results for Comprehensive.
@@
A.8 Multivalue Perturbation
@@
A.9 Statement of Contribution
All co-authors contributed to discussions, provided input on various aspects of the project, and assisted with writing, editing, and advising. In addition to these contributions, FL developed the initial idea, designed and conducted the experiments, contributed significantly to data collection, drafted the paper, and performed all analyses unless otherwise specified. As FL’s mentors during her internship at Microsoft, SM and AW contributed significantly by coordinating resources, guiding the overall direction of the project including data collection, managing the ethical review process, and serving as the primary corresponding authors. On top of advising and paper writing, ELM developed some of the initial experiments for the ablation study and VH contributed to ideation and experiment design.

### v2

@@ 5 Related Work
Previous works on AAVE studies in natural language processing mostly focus on non-reasoning-heavy tasks such as POS tagging (Jørgensen et al., 2015, 2016), language identification and dependency parsing (Blodgett et al., 2016), automatic captioning (Tatman, 2017), and general language generation (Deas et al., 2023). AAVE is also found to be more likely to trigger false positives in hate speech identifiers (Davidson et al., 2019; Sap et al., 2019) due to specific word choices (Harris et al., 2022), be considered negative by automatic sentiment classifier (Groenwold et al., 2020), and cause covert biases in essential areas of social justice (Hofmann et al., 2024). Relevant studies (Ziems et al., 2022; Gupta et al., 2024) also find that rule-based AAVE feature perturbations can downgrade language model performance in a wide range of tasks covered by GLUE (Wang, 2018).
More generally, dialects in world languages pose challenges to natural language processing systems. Ziems et al. (2023) find that auto-encoder models are brittle on rule-based English dialect feature perturbations. Fleisig et al. (2024) report that English dialect speakers perceive responses generated by chatbots to be more negative than Standardized English (SE) prompts. Faisal et al. (2024) find that world dialects cause problems in tasks including dependency parsing (Scherrer et al., 2019) and machine translation (Mirzakhalov, 2021) on mBERT and XLM-R (Conneau et al., 2020).
Fairness and Robustness of Large Language Models. LLMs are widely testified to be both unfair and brittle. They introduce unfair performance (Huang et al., 2023; Dong et al., 2024) and cost (Petrov et al., 2024) to users across different languages, exacerbate social imbalance by marginalizing minority groups in various aspects including gender (Kotek et al., 2023; Fraser and Kiritchenko, 2024), race (Hofmann et al., 2024; Wang et al., 2024), and culture (Naous et al., 2023; Tao et al., 2024). They also show different performance to reasoning tasks in different languages (Huang et al. 2023, 2024; Ranaldi et al. 2024; inter alia). For the first time, our study provides extensive empirical evidence that LLMs exhibit unfairness in reasoning tasks. This bias specifically affects speakers of certain dialects within a single language.
Previous works report that LLMs are very sensitive to slight variations of prompts by introducing typos or paraphrasing in SE (Elazar et al., 2021; Liang et al., 2022; Raj et al., 2022; Zhu et al., 2023b; Lin et al., 2024). In this work, we consider a novel application of using human-written perturbations in AAVE by asking humans to rewrite instances to their dialect and evaluate LLM robustness towards these natural perturbations, which have proven to cause LLMs to be more brittle than synthetic typo-style (Section 4.1) or linguistic-rule-based (Appendix A.9) perturbations.
@@ 6 Conclusion
Our study is the first attempt to objectively evaluate the dialect robustness and fairness of LLMs across reasoning tasks. We introduce ReDial, a dataset comprising over 1.2K parallel prompts in Standardized English and African American Vernacular English (AAVE) tailored to reasoning tasks: algorithm, logic, math, and integrated reasoning. Extensive empirical evidence on ReDial demonstrates that LLMs exhibit significant unfairness and brittleness when reasoning tasks are expressed in AAVE. These findings underscore the unfairness to dialect users and LLMs’ brittleness with natural prompt variations with the same semantics. We advocate for further research to enhance dialect fairness and robustness of LLMs, ensuring equal service for all linguistic groups and demographics.
7 Limitations
First, as the first systematic framework for analyzing LLM bias in dialectal queries for reasoning tasks, we selected AAVE due to its linguistic significance and cultural impact. However, we recognize the vast diversity of dialects worldwide. The insights derived from AAVE may not generalize to other dialects. To ensure annotation quality and maintain the focus of our study, we concentrated on AAVE with high-quality human annotations. Future research could expand on our framework to encompass a wider range of dialects, generating more broadly applicable conclusions.
Second, our benchmark, ReDial, evaluates LLM performance across four categories of reasoning tasks using queries sampled from seven popular and well-documented benchmarks. While these tasks are representative of common reasoning challenges, we acknowledge that reasoning is a multifaceted domain with many additional categories and tasks that fall outside the scope of this study.
Third, we evaluated five representative LLM families in this study, including widely used and sota models. However, given the rapid proliferation of new LLMs, testing every model is infeasible. We hope that future research will use the ReDial benchmark to investigate fairness and reasoning robustness across a broader range of LLMs as they emerge.
Last but not least, while we present extensive empirical evidence demonstrating the performance drop of LLMs on dialectal queries, our study does not deeply investigate the underlying causes of these performance discrepancies or propose systematic methods to mitigate this bias. These topics exceed the scope of our work but are critical for addressing the inequities we have identified. Despite this limitation, we believe that ReDial provides a robust and systematic tool to help researchers explore these issues. The absence of immediate solutions should not detract from the significance of our findings, which lay the groundwork for future efforts to address fairness and robustness in LLMs.
8 Ethic Statement
@@
For data collection, we deliberately do not set hard constraints for annotator identity and demographic verification, recognizing there are no definite boundaries to identify dialects and their speakers (King, 2020). The authors further elaborate that the term “AAVE” itself is contested, with alternatives that could be used instead; in employing the term “AAVE”, we adhere to the widely used terminology in related works on dialects and NLP (Ziems et al., 2022; Gupta et al., 2024). We corroborate the data quality by asking self-identified dialect speakers to cross-validate each others’ answers.
@@ A.1.1 Algorithm
from typing import List def has_close_elements(numbers: List[float], threshold: float) -> bool: """Checkifingivenlistof ␣␣␣␣numbers,areanytwonumbers ␣␣␣␣closertoeachotherthan ␣␣␣␣giventhreshold. ␣␣␣␣$>$has_close_elements([1.0, ␣␣␣␣2.0,3.0],0.5) ␣␣␣␣False ␣␣␣␣$>$has_close_elements([1.0, ␣␣␣␣2.8,3.0,4.0,5.0,2.0],0.3) ␣␣␣␣True ␣␣␣␣""" InstructHumanEval Used in the Paper
@@
Write a python function to remove first and last occurrence of a given character from the string. Your code should pass these tests: assert remove_Occ(“hello",“l") == “heo" assert remove_Occ(“abcda",“a") == “bcd" assert remove_Occ(“PHP",“P") == “H"
@@ A.1.2 Logic
Consider the following premises: “People in this club who perform in school talent shows often attend and are very engaged with school events. People in this club either perform in school talent shows often or are inactive and disinterested community members. People in this club who chaperone high school dances are not students who attend the school. All people in this club who are inactive and disinterested members of their community chaperone high school dances. All young children and teenagers in this club who wish to further their academic careers and educational opportunities are students who attend the school. Bonnie is in this club and she either both attends and is very engaged with school events and is a student who attends the school or is not someone who both attends and is very engaged with school events and is not a student who attends the school." Assuming no other commonsense or world knowledge, is the sentence “Bonnie performs in school talent shows often." necessarily true, necessarily false, or neither? Answer either “necessarily true", “necessarily false", or “neither".
@@ A.4 Data Quality Verification
You will be given two prompts, one in Standard English and one in African American English. Determine whether the African American English prompt is a valid paraphrase of the Standard English prompt. Ignore the semantic validaty of the Standard English prompt. Standard English: "[SAE_PROMPT]" African American English: "[AAVE_PROMPT]" Is the African American English prompt a valid paraphrase of the Standard English prompt?
@@ A.7 Full Results on ReDial
we present the complete results on Redial. Specifically, Table 5 provides the detailed results for Algorithm, Table 6 covers the results for Logic, Table 7 reports the results for Math, and Table 8 reports the results for Integrated Tasks.
@@
Model Setting Original AAVE
[…]
CoT 0.762 0.662(-)0.1
[…]
CoT 0.283 0.058(-)0.225
[…]
CoT 0.075 0.067(-)0.008
[…]
CoT 0.900 0.771(-)0.129
LLaMA-3.1-70B-Instruct Vanilla 0.392 0.113(-)0.279
CoT 0.579 0.500(-)0.079
LLaMA-3-70B-Instruct Vanilla 0.158 0.067(-)0.092
CoT 0.517 0.350(-)0.167
LLaMA-3-8B-Instruct Vanilla 0.025 0.067(+)0.042
CoT 0.029 0.025(-)0.004
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.100 0.075(-)0.025
CoT 0.133 0.071(-)0.062
Mistral-7B-Instruct-v0.3 Vanilla 0.096 0.075(-)0.021
CoT 0.083 0.083(-)-0.0
Phi-3-Medium-128K-Instruct Vanilla 0.050 0.037(-)0.013
CoT 0.067 0.029(-)0.037
Phi-3-Small-128K-Instruct Vanilla 0.058 0.062(+)0.004
CoT 0.096 0.079(-)0.017
Phi-3-Mini-128K-Instruct Vanilla 0.021 0.042(+)0.021
CoT 0.017 0.021(+)0.004
Table 8: All results for Integrated.
A.8 Perplexity of Typos vs. Natural AAVE Text
We gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop did not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. Interestingly, larger models such as LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct degrade less on synthetic noise yet still struggle disproportionately with authentic AAVE data. This suggests their brittleness stems from reliance on SE priors and limited encoding of AAVE-specific discourse structures.
Figure 4: Model performance on misspelled SE compared to human-written AAVE data. We gradually add noise to SE ReDial to increase its perplexities until they surpass the perplexity of AAVE ReDial and report the models’ performance on every perturbation level. Horizontal and vertical lines refer to model pass rates/perplexities on AAVE ReDial respectively. Larger LLMs (i.e., LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct) perform worse on AAVE than on perturbed text with a similar perplexity level.
A.9 Multivalue Perturbation
@@
A.10 Qualitative Analysis of Standardized Prompting
We observe a sensible gap between model performance on Standardized ReDial with zero-shot prompting and AAVE ReDial with standardization prompting. In this section, we qualitatively compare GPT-4o’s outputs in these two settings, the model with the best absolute overall performance, and examine its errors. We focus on the math subset of ReDial and identify three key error patterns: wrong question rephrasing, distraction by irrelevant information, and failure to execute all steps.
A.11 Synthetic Lexical Feature list
Following Ziems et al. (2022), we implement a feature list with distinct AAVE lexicons: [’got’, ’ain’t’, ’no’, "’", ’gonna’, ’wanna’, ’gotta’, ’done’, ’been’, ’finna’, ’gunna’, ’gon’,], and compute the Spearman’s correlation between their frequencies in AAVE ReDial and the performance drop.
A.12 Statement of Contribution
All co-authors contributed to discussions, provided input on various aspects of the project, and assisted with writing, editing, and advising. In addition to these contributions, FL developed the initial idea, designed and conducted the experiments, contributed significantly to data collection, drafted the paper, and performed all analyses unless otherwise specified. As FL’s mentors during her internship at Microsoft, SM and AW contributed significantly by coordinating resources, guiding the overall direction of the project including data collection, managing the ethical review process, and serving as the primary corresponding authors. XW conducted experiments on Claude and GPT-o1. On top of advising and paper writing, ELM developed some of the initial experiments for the ablation study and VH contributed to ideation and experiment design.